OpenStudy (anonymous):

please help :( classify the system as an independant, dependant, or inconsistant. best answer and i will become a fan. problem in comments

OpenStudy (anonymous):

\[2x-4y=22\]\[5x-10y=16\]

OpenStudy (anonymous):

first lets divide the first equation by 2 so we get x-2y=11 now lets multiply by -5 we get -5x+10y=-55 lets sum with 5x-10y=16 we get 0=-39 so it's inconsistent

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem